OpenStudy (anonymous):

Tony has 14 fewer than twice as many hits at Kal. Nick had half as many hits as Tony. Combined they have 371 hits. How many hits did each of them have. Solve using an algebraic method

OpenStudy (agent47):

Kal: h Tony: 2h-14 Nick: (1/2)(2h-14) Kal+Tony+Nick = 371 h+(2h-14)+(1/2)(2h-14)=371 h+2h-14+h-7=371 4h=371+21 4h=392 h=98 <- Kal has 98 hits

OpenStudy (agent47):

Tony has: 98*2-14=182

OpenStudy (agent47):

Nick has half as many as Tony: 182/2=91

OpenStudy (anonymous):

thanks so much